Circular No. 8657

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005nb, 2006B, 2006C Further to IAUC 8622, R. Quimby, P. Mondol, and F. Castro, Univer-

sity of Texas, report the discovery of a supernova in unfiltered CCD images taken on 2005 Dec. 17.38, 2006 Jan. 3.34, and 7.35 UT (each showing the new object at mag ∼ 17.2) using the 0.45-m ROTSE-IIIb telescope at the McDonald Observatory. SN 2005nb is located at α = 12^{h}13^{m}37^{s}.61, δ = +16^{o}07′16′′.2 (equinox 2000.0; uncertainty ± 0′′.9), which is 1′′.5 west and 5′′ north of the center of UGC 7230 (the brighter galaxy in the pair). The supernova was found by subtracting a co-addition of images taken between 2005 Jan. 7 and Apr. 12 (limiting mag ∼ 18.8). A spectrum (range 420– 890 nm) of 2005nb, obtained on Jan. 8.39 UT with the 9.2-m Hobby-Eberly Telescope (+ Marcario Low-Resolution Spectrograph) by B. Roman and S. Rostopchin, shows it to be a type-Ic supernova (similar to the spectrum of SN 2002ap at 10 days past maximum (Gal-Yam et al. 2002, via the SUS-

PECT archive at http://bruford.nhn.ou.edu/~suspect/index1.html).

Two more apparent supernovae have been found on unfiltered CCD images: 2006B by T. Puckett and G. Sostero (cf. IAUC 8655; 0.60-m automated supernova patrol frames); and 2006C by W. Li (LOSS/KAIT; cf. IAUC 8647). Additional magnitudes for 2006B in UGC 12538: 2005 Sept. 10, [19.3 (Puckett); 2006 Jan. 8.02, 16.5 (Sostero, 0.45-m reflector near Udine, Italy). SN 2006B is a type-Ia supernova (cf. CBET 360). Additional magnitudes for 2006C: 2004 May 25.31, [19.5; 2006 Jan. 9.53, 17.4.

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006B Jan. 8.02 23^{h}21^{m}09.^{s}78 +33^{o}24′01′′.2 16.7 1′′.0 W, 2′′.9 N
2006C Jan. 6.54 12 02 32.50 +41 02 40.2 17.5 45′′.3 W, 34′′.2 S

COMETS C/2005 W14 AND C/2005 W15 (SOHO) Additional Kreutz comets have been found on SOHO website images

(cf. IAUC 8654). C/2005 W14 was small, dim, and stellar in appearance on C3 images, while on C2 images it appeared slightly diffuse, reaching mag 7.1 at 6.0R_{⊙} on Nov. 23.838. C/2005 W15 was small, faint, and stellar in appearance on C3 images, and very faint and diffuse on C2 images.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 W14 Nov. 23.571 15^{h}53.^{m}0 −22^{o}57′ C3/2 TS 2005-Y22
C/2005 W15  25.029 16 01.3 −24 08 C3/2 HS 2005-Y22
2006 January 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green